the square of an integer is 30 more than the integer. find the integer.
Answer by nerdybill(7384)

the square of an integer is 30 more than the integer. find the integer.
.
Let x = the integer
then
x^2 = x + 30
x^2 -x -30 = 0
(x-6)(x+5) = 0
x = {-5,6}
.
Therefore, the integer could either be -5 or 6